OpenStudy (anonymous):

1/y+5 + 3/5y can someone help me

OpenStudy (anonymous):

to add them, you need the same denominator. Therefore, we multiply 1/y+5 by 5y and 3/5y by y+5 So we get: (6y+5)/(5y^2+25y)

OpenStudy (anonymous):

\[\frac{1}{y+5}+\frac{3}{5 y}=\frac{8 y+15}{5 y^2+25 y} \] Checks out with y evaluated at 31.